OpenStudy (anonymous):

(45,0), (0, 1.39), and (18, 3.2) Use those to find the equation of the quadratic. x=ax^2+bx+c

OpenStudy (mathmale):

This is quite an interesting problem (at least to me). Each of the 3 points contains info on x and y: In (45,0), x=45 and y=0. Substitute those values into the quadratic y=ax^2+bx+c. You'll come up with an equation in a, b and c. Do the same thing for the other two points. You'll end up with 3 equations in a, b and c. Your job is to solve this system of linear equations for a, b and c. Then write out the original quadratic, substituting your values of a, b and c.

OpenStudy (anonymous):

Whats The Original Equation Look Like?

OpenStudy (mathmale):

The "original equation" is your x=ax^2+bx+c. all you need to change here is that initial x: You meant y=ax^2+bx+c.

OpenStudy (anonymous):

Oh Okay Thank You And May I Ask If These Look Alright: 2025a+45b+c c=1.39 3.2=324a+18b+c

OpenStudy (mathmale):

@LannaBoo24: The first expression you've typed in is not an equation. Can you fix this, so that you DO have an equation there?

OpenStudy (anonymous):

Un Square It?

OpenStudy (mathmale):

Notice that you don't have an ' = ' sign in the first expression you've typed in. That means you have an expression, not an equation. What you are trying to say in writing this equation is that "when x=45, y=0." Can you now fix the first expression (make it an equation)?

OpenStudy (mathmale):

2025a+45b+c=0. Now you have 3 equations in the three unknowns, a, b and c. Please solve for a, b and c.

OpenStudy (anonymous):

C=1.39 So I Can Just Put That Into The Other Two Equations Like This: 2025a+45b+1.39=0 324a+18b+1.39=3.2 Right?

OpenStudy (mathmale):

Sure! That's one of the faster ways of solving this problem. Now you need only find the values of a and b.

OpenStudy (anonymous):

Okkie Dokkie c: Well On The 2nd Equation That Has To Be Equal To 0 Also Or Is It Okay Like It Is?

OpenStudy (mathmale):

Actually, your second equation is 0a + 0b +1c = 1.39, so your c=1.39 is correct.

OpenStudy (anonymous):

Opps. I Meant In The 324a+18b+1.39=3.2 Sorry.

OpenStudy (mathmale):

That you can find your own mistakes shows that you understand a great deal of what is happening here. :)

OpenStudy (mathmale):

You're on the right track. I need to get off the Internet now, unfortunately. Good luck!!

OpenStudy (anonymous):

Thank You c: